Mary Arnold – Faith and the Everyday Practice of Science

Mary, this is a very good start on an essay concerning Grinnell’s ideas about complementarity of science and religion. However, I don’t think you have really “nailed” a description of the differences between the two fields, at least according to common thinking (what Grinnell is arguing against). Your first paragraph points to the distinction between pragmatic concerns and theoretical concerns. But that is not so clear-cut (see comment 1 below). Your second paragraph says the fundamental conflict between science and religion is in how they draw conclusions. There certainly is a difference there, but is that the most central difference? It might be good to start a revised version of this (if you decide to write one) by addressing another question: What are the ultimate goals of science and religion, and how are they different (and/or similar)?

Specific Comments

1. It sounds as though you are equating science with the pragmatic and religious with the theoretical here. But I would say that science and mathematics certainly have a theoretical side as well as a pragmatic side. So the difference between science and religion is not only the difference between pragmatic and theoretical concerns.

2. Everyday Practice of Science is a rather unclassifiable book, so it could be called “novel,” meaning new or unusual. I’m guessing that is what you were thinking here. But when I read this the first time, it seemed that you were calling the book a work of fiction. It would be better to rewrite this to eliminate the possible confusion.

3. I think it would be better to say here that courts held that the introduction of creationist science curricula was a breach of the separation of church and state. This is a somewhat controversial legal issue. Your flat statement makes it sound too cut and dried. Also, it’s interesting to consider the way creationism is often presented as a “scientific” theory and an alternative to Darwinian evolution. Why do you suppose the supporters of creationism try to do that?

4. This is still a part of an expression of Grinnell’s opinion. It would be good to make that clearer because there are certainly lots of scientists who would disagree violently with this statement.

5. I am not sure what you meant by this. Do you mean science cannot explain the human feelings of a connection to the divine??

6. This idea that embracing complementarity could lead to mutual respect and acceptance between religious and scientific viewpoints is a very good one. You could say more about that!

Amanda Cabrera – One Natural World, Two Visions

I got your email message accompanying the paper, and I agree that this is still somewhat rough and unfinished. I hope you will have the time to think over the following comments and resubmit a revised version of the paper.

Specific Comments

1. Saying “integrated” here sounds a little strange. Many scientists might say that when religious concerns impact how science is taught and practiced, then it is more an intrusion than an integration. (The primary meaning of integration is inclusion or mixing in a more neutral way.) Your second sentence is making a very different point (following Grinnell). So this probably deserves to be expanded and explained a bit more – you need to give the reader a clue that you will be dealing with two very different things.

2. The people who have been agitating for “creationist science” in schools are primarily fundamentalist Protestants. In recent times at least, the Catholic Church has been much less prone to that sort of literal interpretation of Scripture and much more open to modern geological theories of the age of the Earth and biological theories of evolution. So I think this sentence is somewhat inaccurate.

3. Grinnell’s idea of “intelligible design” is also a very good point to spend more time on. The idea here is the feeling (which comes down to a form of faith, for Grinnell at least) that there are underlying patterns in Nature that humans are able to understand.

4. I don’t really see the connection between what you are saying about Biblical literalism and Grinnell’s complementarity here. This is something you probably want to rethink.

5. Is it faith that is expressed through the reasoning of where to go on the bicycle? Isn’t this sort of reasoning more closely related to ethics or morality? Aren’t those different from faith? (That is, religion also includes ethical and moral statements of what we should do.)

6. Could you say more about why you think Grinnell’s dabbling in philosophy is dangerous here? Are you saying he is not qualified to offer those opinions because he is “only” a scientist?

Stephanie Craig – The Battle Between Effect and Passion

I like your first draft paper very much – it has a number of good points and it strikes a good balance between your personal experience and the ideas expressed by Grinnell in his book. There are a few places where you had typos or minor slips. Also there were a few examples of what I would call overly colloquial language – “zapping the life out of” (page 1) in particular.

Specific Comments

1. I gather that a lot of students feel the way you describe here (that is, that having to explain a solution takes the fun of out finding the solution). Could you say more about why you have that feeling? I don’t really understand that myself, since I always think that the best way to make sure you really understand something is to try to explain it to someone else. Any gaps in one’s thinking become painfully clear then, either because of the uncomprehending stares that you get from one’s audience, or from the feeling of “hand-waving” that comes from trying to dance around a point that is not entirely nailed down.

2. I like this example very much – I think it is exactly right!

3. You are also correct that statistics is often used in the way you describe – as a way to summarize a large collection of information and make it easier to deal with. You seem to be saying that this is like the process of writing a short and to the point research article without any of the distracting extra information that might have been developed along the way to the final results. Am I catching your point here, or did you mean something different? Also, aren’t there times when you might discover new things in the process (patterns that were not apparent in the whole data set at the start). That is, couldn’t working with the statistics be part of the discovery process, in addition to or rather than the credibility process?

Josh Hauser – Practice of Science and Mathematics

Your first paragraph was a little rough (see the suggested corrections marked on the paper). But after you got started, your writing improved quite a bit. One suggestion – “kids” sounds too informal for this kind of writing. What you really mean is “students” so say it that way.

I agree with a lot of what you are saying here, and you present your points pretty well.

Specific Comments

1. This is a very important point. This kind of training in logical thinking and mental discipline is also one of the traditional reasons for including mathematics in school curricula. This goes all the way back to the ancient Greeks.

2. I’m not sure it’s really accurate to say that there is “much room for extra information” in research articles. After all, textbooks are usually hundreds of pages long, but articles are only rarely longer than 20 or 30 pages. There is a difference too in the level of knowledge assumed of the readers. In research articles, the audience is the community of experts in a field, so not all of the background needs to be filled in.

3. I’m not sure I follow what you are saying here. It is not that common, I think, for later articles to re-work the data from earlier ones. So I don’t think “taking the results from a previous experiment in order to find different patterns and characteristics” happens that often (at least not in the parts of science I know the most about). Instead, people might try to reproduce other scientists’ results by redoing the experiments in a different way and generating new data of their own.

Paige Carr – Faith in Religion and Science of the Natural World

Your essay is well-structured and presents a good synthesis of Grinnell’s ideas about the relation between faith, science and religion. I only have a few small corrections and comments.

Specific Comments

1. You say that “without religion, there would be no point in discovering the reason why things in life work.” Does Grinnell say anything like that? Or is this a personal opinion of your own? It would be good to say more about this and try to justify it, because I think this is a statement that many people might disagree with. Isn’t it possible for an atheist to be a scientist and want to understand how nature works? Can’t an atheist be a good, moral person and be concerned with the welfare of others? Is religion the only source of moral and ethical values that provide the basis for a good life?

2. Toward the end of the essay when you are discussing Grinnell’s idea of complementarity, there are a few instances where you are misusing the word. This is a noun – the property of being complementary. The adjective form is complementary.